OpenStudy (anonymous):

What value in place of the question mark makes the polynomial below a perfect square trinomial? 9x2+?x+64

OpenStudy (anonymous):

you need so square \[(3x+8)^2\] and see what the middle term would be

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

Yeah, it's 48

OpenStudy (anonymous):

\[(3x+8)(3x+8)=9x^2+3x\times 8+ 3x\times 8 + 64\] \[=3x^2+48x+64\] you got it